OpenStudy (anonymous):

4. Factor the following trinomials: x2 + 8x + 15

OpenStudy (anonymous):

i think using complete the square is the easiest for this one

myininaya (myininaya):

\[(x+3)(x+5)\]

myininaya (myininaya):

since 3(5)=15 and 3+5=8

OpenStudy (anonymous):

x^2+8x+15 \[(x^2+8x+16) -1\] \[(x^2+4x+4x+16)-1\] \[(x+4)+4(x+4) -1\] \[(x+4)(x+4) - 1\] \[(x+4)^2-1\] \[(x+4-1)(x+4+1) \] since a^2-b^2=(a+b)(a-b)